Compute each of the double double sums below

(a)3∑(i=1) 2∑(j=2) (i+j)

(b)3∑(i=0) 2∑(j=0) (2i+3j)


1
Expert's answer
2022-05-19T17:54:51-0400

(a)


"\sum_{i=1}^{3}\sum_{j=2}^{2}( i+j) =\sum_{i=1}^{3}(i+2)""=1+2+2+2+3+2=12"

(b)


"\sum_{i=0}^{3} \sum_{j=0}^{2} (2i+3j)""= \sum_{i=0}^{3}( 2i+(2i+3)+(2i+6))""= \sum_{i=0}^{3}( 6i+9)""=9+6+9+6\cdot2+9+6\cdot3+9"

"=72"